OpenStudy (anonymous):

ram has already given five math test. if he scores 70 on upcoming sixth exam, his average mark goes down by 4 points. what is his average score now?

OpenStudy (ajprincess):

Let the average of five subjects be x. Then the total marks scored in five tests=5*x Total marks scored in 6 tests=5x+70 Average=Total marks/number of tests x-4 =(5x+70)/6 6(x-4)=5x+70 Solve for x. Then to find the current average subtract 4 from the value of x. Can u do that @shreejanramdev?